OpenStudy (anonymous):

What is the slope of AD help i put A am i right? (Points : 2) -1 1 -4/3 -2/5

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Directrix I PUT A AM I RIGHT?

Directrix (directrix):

I did not get 1 for the slope. Did you use (-5,2) and (-3,4) as the coordinates of D and A, respectively? I think there may be a sign error in the answer. Check your work and I'll check mine.

OpenStudy (anonymous):

K

Directrix (directrix):

I think I messed up. The line determined by the points has to have a positve slope because the line is rising. m = [ (2 - 4)] /[ (-5) - (-3) ] = ? What does m equal?

OpenStudy (anonymous):

-2/8=-4

Directrix (directrix):

What does (2 - 4) =

jimthompson5910 (jim_thompson5910):

visually you can eliminate choice A why? because AD is a line that slopes upward as you move from left to right so the slope must be some positive number

OpenStudy (anonymous):

-2

Directrix (directrix):

Also, what does (-5) - (-3) =

OpenStudy (anonymous):

8

Directrix (directrix):

Right on -2 but not the second calculation. (-5) - (-3) = -5 + 3 = ?

Directrix (directrix):

@tigersandbulls Waiting for you on answer to -->(-5) - (-3) = -5 + 3 = ?

OpenStudy (anonymous):

8=-2 @Directrix SRRY I WAS TAKING MY DOG OUT

Directrix (directrix):

So, the slope m = [ (2 - 4)] /[ (-5) - (-3) ] = (-2)/(-2) = ??? @tigersandbulls

OpenStudy (anonymous):

-2/8=1

OpenStudy (anonymous):

ITS D

Directrix (directrix):

Where did you get the 8? D is not the answer. The slope cranks out to (-2)/(-2). What is -2 divided by -2 ? -----> @tigersandbulls

Directrix (directrix):

@tigersandbulls